# Integer quadratic forms and extensions of subsets of linearly independent roots

### v2

@@ 1.2. Root system, Cartan matrix, Carter diagram
Let $\varPhi$ be a finite root system, $W$ be the corresponding finite Weyl group. We consider only root systems with simply-laced Dynkin diagrams. The Cartan matrix associated with $\varPhi$ is denoted by $B$. Let $\mathscr{B}$ be the quadratic Tits form associated with $B$ and (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) be the inner product induced by $\mathscr{B}$. For the Cartan matrix $B$, the following well-known property [^1] For the matrix $B/2$, eq. (1.3) looks like this: $\mathscr{B}(\alpha)=1\Longleftrightarrow\alpha\in\varPhi$, see [K80, Prop. 1.6]. holds:
@@
$\begin{split}&(a)\text{ The nodes of }\Gamma\text{ correspond to a subset of linearly independent roots in }\varPhi.\\ &(b)\text{ If a subdiagram of }\Gamma\text{ is a cycle, then it contains an even number of nodes.}\end{split}$ (1.4)
@@ 1.3. The partial Cartan matrix
$(\tau_{i},\tau_{j})=\begin{cases}0,&\text{if }\tau_{i}\text{ and }\tau_{j}\text{ are not connected},\\ -1,&\text{if edge }\{\tau_{i},\tau_{j}\}\text{ is solid},\\ 1,&\text{if edge }\{\tau_{i},\tau_{j}\}\text{ is dotted}.\\ \end{cases}$
@@ 1.5. Homogeneous Carter diagrams and the transition matrix
$\begin{array}[]{ll}&(1)\;\{D_{4}(a_{1}),D_{4}\}\\ &(2)\;\{D_{l}(a_{k}),D_{l}\}\\ &(3)\;\{E_{6}(a_{1}),E_{6}\}\\ &(4)\;\{E_{6}(a_{2}),E_{6}(a_{1})\}\\ &(5)\;\{E_{7}(a_{1}),E_{7}\}\\ &(6)\;\{E_{7}(a_{2}),E_{7}\}\\ &(7)\;\{E_{7}(a_{3}),E_{7}(a_{1})\}\\ &(8)\;\{E_{7}(a_{4}),E_{7}(a_{3})\}\\ \end{array}\qquad\qquad\begin{array}[]{ll}&~{}(9)\;\{E_{8}(a_{1}),E_{8}\}\\ &(10)\;\{E_{8}(a_{2}),E_{8}\}\\ &(11)\;\{E_{8}(a_{3}),E_{8}(a_{2})\}\\ &(12)\;\{E_{8}(a_{4}),E_{8}(a_{1})\}\\ &(13)\;\{E_{8}(a_{5}),E_{8}(a_{4})\}\\ &(14)\;\{E_{8}(a_{6}),E_{8}(a_{4})\}\\ &(15)\;\{E_{8}(a_{7}),E_{8}(a_{5})\}\\ &(16)\;\{E_{8}(a_{8}),E_{8}(a_{7})\}\\ \end{array}$ (1.8)
@@ 1.6. Transitions
$\begin{cases}M\tau_{i}=\tau_{i}\text{ for all }\tau_{i}\in\widetilde{S},\tau_{i}\neq\widetilde{\alpha},\\ M\widetilde{\alpha}=\alpha=-\widetilde{\alpha}+\sum t_{i}\tau_{i},\text{ the sum is taken over all }\tau_{i}\in\widetilde{S}\text{ except }\widetilde{\alpha};~{}t_{i}\in\mathbb{Z}.\\ \end{cases}$ (1.10)
@@
$M\widetilde{\alpha}=\alpha\text{ and }M\alpha=\widetilde{\alpha}.$
@@ 1.7. Admissible diagrams and Carter diagrams
According to (1.4$(a)$), there exists a certain set (1.12) of linearly independent roots.Thanks to (1.4$(b)$), there exists a partition $S=S_{\alpha}\coprod S_{\beta}$ which is said to be the bicolored partition. The admissible diagram is bicolored, i.e., the set of nodes can be partitioned into two disjoint subsets $S_{\alpha}=\{\alpha_{1},\dots,\alpha_{k}\}$ and $S_{\beta}=\{\beta_{1},\dots,\beta_{h}\}$, where roots of $S_{\alpha}$ (resp. $S_{\beta}$) are mutually orthogonal.
@@ 2.1. When a vector is a linkage root?
[…]
@@ 2.2. Criterion that a vector is a linkage root
${}^{t}{M}\cdot B_{\Gamma^{\prime}}\cdot M=B_{\Gamma}.\$
@@ 2.3. Full linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$
where $\varPhi(\widetilde{\Gamma})$ is the root system associated with $\widetilde{\Gamma}$, $S^{\prime}$ some $\Gamma^{\prime}$-set and $L(S^{\prime})$ is the linear span ${\rm Span}(S^{\prime})$. It was shown in Proposition 4.2 that $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime})$ does not depend on choosing $\Gamma^{\prime}$-set $S^{\prime}$. Then, the set (2.4) can be denoted by $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime})$. The set $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime})$ is said to be partial linkage system $\widetilde{\Gamma}$ over $\Gamma^{\prime}$. The pair $\{\Gamma^{\prime},\widetilde{\Gamma}\}$, where $\Gamma^{\prime}\in C(\Gamma)$, is said to be the vertex extension and is denoted by $\Gamma^{\prime}\prec\widetilde{\Gamma}$. The union of all partial systems by all possible vertex extensions is said to be the linkage system or full linkage system of $\Gamma$:
@@
$D_{4}$,  $D_{4}(a_{1})$ $3$ $3\times 8=24$ $24$
$D_{5}$,  $D_{5}(a_{1})$ $3$ $10$ $2\times 16=32$ $42$
$D_{6}$, $D_{6}(a_{1})$, $D_{6}(a_{2})$ $3$ $12$ $2\times 32=64$ $76$
$D_{7}$, $D_{7}(a_{1})$, $D_{7}(a_{2})$ $3$ $14$ $2\times 64=128$ $142$
$D_{l}$,  $D_{l}(a_{k})$, $l>7$ $1$ $2l$ $2l$
@@ 3.1.1. Action of $B^{-1}_{\Gamma}$ on the space $L$
Let $S=\{\tau_{1},\dots,\tau_{l}\}$ be a $\Gamma$-set, $L$ the space spanned by roots of $S$, $\gamma$ be a root in $L$:
@@ 3.1.2. The projection of the linkage root
Let $L^{\perp}$ be the orthogonal complement of $L$ to $L(\gamma)$ in the sense of the symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) associated with the root system $\varPhi$:
@@
(ii) The vector $\mu$ is, up to sign, a fixed vector for any root $\theta\in L(\gamma)\backslash L$.
(iii) The value $\mathscr{B}_{\Gamma}(\theta_{L})$ is constant for any root $\theta\in L(\gamma)\backslash L$.
(iv) The vector $\theta=\theta_{L}+\mu$ is a root if and only if the conjugate vector $\overline{\theta}=\theta_{L}-\mu$ is a root. In addition, $\theta\in L(\gamma)\backslash L$ if and only if $\overline{\theta}\in L(\overline{\gamma})\backslash L$.
@@
(vi) If $\delta$ is a root in $L(\gamma)\backslash L$ such that $\delta^{\nabla}=\gamma^{\nabla}$, then $\delta=\gamma$ or $\delta=\overline{\gamma}$.
(vii) For any root $\theta\in L(\gamma)\backslash L$, we have
@@
and, $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})$ is a constant for all roots $\theta\in L(\gamma)\backslash L$.
@@
$\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})=\langle B_{\Gamma}^{-1}\theta^{\nabla},\theta^{\nabla}\rangle=\langle{\theta_{L}},B_{\Gamma}\theta_{L}\rangle=\mathscr{B}_{\Gamma}(\theta_{L}).\$
@@ 3.3. Quadratic forms for homogeneous Carter diagrams
${}^{t}{M}\cdot B_{\Gamma^{\prime}}\cdot M=B_{\Gamma},\$ (3.8)
@@ 3.4. The chain of homogeneous pairs
$\{\{\Gamma_{0},\Gamma_{1}\},\{\Gamma_{1},\Gamma_{2}\},\dots,\{\Gamma_{k-1},\Gamma_{k}\},\text{ where }\Gamma_{0}=\Gamma^{\prime}\text{ and }\Gamma_{k}=\Gamma.\$ (3.9)
@@
$\begin{cases}M^{(j)}(\tau^{j}_{i})=\tau^{j+1}_{i}\text{ for some }i\in\{1,\dots,n\},\\ M^{(j)}(\tau^{j}_{k})=\tau^{j}_{k}\quad\text{ for }k\neq i.\\ \end{cases}\$
@@ 3.5. Spectrum of the Dynkin and Carter diagram
Proof. Note that for any transition matrix $M$ and any similarity matrix $L_{\tau}$ the determinant is equal to $-1$, see (1.10). Then, by (3.8) and (3.10), we have
@@
$42\cos\varphi+2\;=\;(\rho-2)^{2},\;\text{ i.e., }\;(\rho-2)^{2}\;=\;4\cos^{2}\frac{\varphi}{2}\;\leq\;$ 4 .
Therefore, $0\leq\rho\leq 4$. By Proposition 3.1(ii) $B_{\Gamma}$ is positive definite, then $\rho>0$. For the finite Weyl groups, $1$ is not an eigenvalue of the Coxeter transformation, then $\rho\neq 4$. Thus, $40<\rho<$ 4. ∎
@@ 3.6.2. Action of the Weyl group $W_{S}^{\vee}$
$(s^{*}_{\tau_{i}}u^{\nabla})_{\tau_{k}}=\begin{cases}-u^{\nabla}_{\tau_{i}},&\text{ if }k=i,\\ u^{\nabla}_{\tau_{k}}+u^{\nabla}_{\tau_{i}},&\text{ if }\{\tau_{k},\tau_{i}\}\text{ is a {\it solid} edge, i.e., }(\tau_{k},\tau_{i})=-1,\\ u^{\nabla}_{\tau_{k}}-u^{\nabla}_{\tau_{i}},&\text{ if }\{\tau_{k},\tau_{i}\}\text{ is a {\it dotted} edge, i.e., }(\tau_{k},\tau_{i})=1,\\ u^{\nabla}_{\tau_{i}},&\text{ if }\tau_{k}\text{ and }\tau_{i}\text{ are not connected, i.e., }(\tau_{k},\tau_{i})=0.\end{cases}$ (3.20)
@@
Proof. Let $\{\tau_{k},\tau_{i}\}$ be a solid edge. If $u^{\nabla}_{\tau_{k}}=u^{\nabla}_{\tau_{i}}=1$ (resp. $u^{\nabla}_{\tau_{k}}=u^{\nabla}_{\tau_{i}}=-1$) then roots $\{-u,\tau_{k},\tau_{i}\}$ (resp. $\{u,\tau_{k},\tau_{i}\}$) constitute the root system corresponding to the extended Dynkin diagram $\widetilde{A}_{2}$, which is impossible, see [St17, Lemma A.1]. For remaining pairs $\{u^{\nabla}_{\tau_{k}},u^{\nabla}_{\tau_{i}}\}$, we have $-1\leq u^{\nabla}_{\tau_{k}}+u^{\nabla}_{\tau_{i}}\leq 1$. Now, let $\{\tau_{k},\tau_{i}\}$ be a dotted edge. If $u^{\nabla}_{\tau_{k}}=1$ and $u^{\nabla}_{\tau_{i}}=-1$ (resp. $u^{\nabla}_{\tau_{k}}=-1$ and $u^{\nabla}_{\tau_{i}}=1$) then roots $\{u,-\tau_{k},\tau_{i}\}$ (resp. $\{u,\tau_{k},-\tau_{i}\}$) constitute the root system $\widetilde{A}_{2}$, which is impossible. For remaining pairs $\{u^{\nabla}_{\tau_{k}},u^{\nabla}_{\tau_{i}}\}$, we have $-1\leq u^{\nabla}_{\tau_{k}}-u^{\nabla}_{\tau_{i}}\leq 1$. ∎
@@
$\begin{split}(s_{\tau_{i}}\gamma)^{\nabla}=&\left(\begin{array}[]{c}(s_{\tau_{i}}\gamma,\tau_{1})\\ \dots\\ (s_{\tau_{i}}\gamma,\tau_{l})\\ \end{array}\right)=\left(\begin{array}[]{c}(s_{\tau_{i}}\gamma_{L}+\mu,\tau_{1})\\ \dots\\ (s_{\tau_{i}}\gamma_{L}+\mu,\tau_{l})\\ \end{array}\right)=\left(\begin{array}[]{c}(s_{\tau_{i}}\gamma_{L},\tau_{1})\\ \dots\\ (s_{\tau_{i}}\gamma_{L},\tau_{l})\\ \end{array}\right)=(s_{\tau_{i}}\gamma_{L})^{\nabla}.\\ \end{split}$
@@
$\begin{split}&\mathscr{B}_{\Gamma}(s_{\tau_{i}}{v})=\langle B_{\Gamma}s_{\tau_{i}}{v},s_{\tau_{i}}{v}\rangle=\langle s^{*}_{\tau_{i}}{B}_{\Gamma}{v},s_{\tau_{i}}{v}\rangle=\langle{B}_{\Gamma}{v},{v}\rangle=\mathscr{B}_{\Gamma}(v).\\ \hskip 65.44142pt&\mathscr{B}_{\Gamma}^{\vee}(s^{*}_{\tau_{i}}{u^{\nabla}})=\langle B_{\Gamma}^{\vee}s^{*}_{\tau_{i}}{u^{\nabla}},s^{*}_{\tau_{i}}{u^{\nabla}}\rangle=\langle s_{\tau_{i}}{B}_{\Gamma}^{\vee}{u^{\nabla}},s^{*}_{\tau_{i}}{u^{\nabla}}\rangle=\langle{B}_{\Gamma}^{\vee}{u^{\nabla}},{u^{\nabla}}\rangle=\mathscr{B}_{\Gamma}^{\vee}(u^{\nabla}).\hskip 65.44142pt\\ \end{split}$
@@ 4.1. The vertex extension
In this chapter, we will use definition $L:=L(S)=L(S^{\prime})$. Let $\widetilde{\Gamma}$ be a Dynkin diagram with the root system $\varPhi(\widetilde{\Gamma})$ such that
$\begin{split}(1)&\quad{\rm rank}(\varPhi(\widetilde{\Gamma}))={\rm rank}(\varPhi(\Gamma))+1,\\ (2)&\quad L\subset{\rm Span}(\varPhi(\widetilde{\Gamma})).\end{split}$
@@
$\begin{array}[]{c}B_{\Gamma}=\left[\begin{array}[]{ccccc}2&0&0&-1&0\\ 0&2&0&-1&-1\\ 0&0&2&-1&0\\ -1&-1&-1&2&0\\ 0&-1&0&0&2\\ \end{array}\right]\begin{array}[]{c}\alpha_{1}\\ \alpha_{2}\\ \alpha_{3}\\ \beta_{1}\\ \beta_{2}\\ \end{array}\end{array},\qquad B_{\Gamma}^{-1}=\begin{array}[]{c}\frac{1}{4}\left[\begin{array}[]{ccccc}5&4&3&6&2\\ 4&8&4&8&4\\ 3&4&5&6&2\\ 6&8&6&12&4\\ 2&4&2&4&4\\ \end{array}\right]\begin{array}[]{c}\alpha_{1}\\ \alpha_{2}\\ \alpha_{3}\\ \beta_{1}\\ \beta_{2}\\ \end{array}\end{array}$
@@ 4.2. Partial linkage systems
where $\varPhi(\widetilde{\Gamma})$ is the root system associated with $\widetilde{\Gamma}$, $S^{\prime}$ some $\Gamma^{\prime}$-set and the linear space $L=L(S^{\prime})$ is given by (4.1).
@@
Proof. Let $S^{\prime}_{1}$ and $S^{\prime}_{2}$ be two different $\Gamma^{\prime}$-sets in $\varPhi(\widetilde{\Gamma})$, and let $S_{i}^{\prime}=\{\tau_{1}^{i},\dots,\tau_{l}^{i}\}$ for $i=1,2$. Consider vector $\gamma_{1}\in\varPhi(\widetilde{\Gamma})$ obtained by the vertex extension of $\Gamma^{\prime}$ using the $\Gamma^{\prime}$-set $S^{\prime}_{1}$. Let $\gamma_{1},L$ be the projection of $\gamma$ onto $L=L(S_{1}^{\prime})$. By (3.3) we have
$\gamma_{1}=\mu_{1}+\sum\limits_{j=1}^{l}t_{j}\tau^{1}_{j},\$ (4.3)
where $\{t_{j}\mid j=1\dots l\}$ are coefficients of decomposition of $\gamma_{1},L$ by the basis $\{\tau_{1}^{i},\dots,\tau_{l}^{i}\}$. From $\gamma_{1}$ we construct $\gamma_{2}\in\varPhi(\widetilde{\Gamma})$ by the same vertex extension using $\Gamma^{\prime}$-set $S^{\prime}_{2}$:
$\gamma_{2}=\mu_{2}+\sum\limits_{j=1}^{l}t_{j}\tau^{2}_{j},\$ (4.4)
@@
Since inner products $(\tau^{1}_{j},\tau^{1}_{k})$ and $(\tau^{1}_{j},\tau^{1}_{k})$ correspond to the same edge of $\Gamma^{\prime}$ (or both are equal to 00) we get $\gamma^{\nabla}_{1}=\gamma^{\nabla}_{2}$. Thus, we get

### v3

@@ 1.2. Root system, Cartan matrix, Carter diagram
Let $\varPhi$ be a finite root system, $W$ be the corresponding finite Weyl group. We consider only root systems with simply-laced Dynkin diagrams. The Cartan matrix associated with $\varPhi$ is denoted by $B$. Let $\mathscr{B}$ be the quadratic Tits form associated with $B$ and $(\cdot,\cdot)$ be the inner product induced by $\mathscr{B}$. For the Cartan matrix $B$, the following well-known property [^1] For the matrix $B/2$, eq. (1.3) looks like this: $\mathscr{B}(\alpha)=1\Longleftrightarrow\alpha\in\varPhi$, see [K80, Prop. 1.6]. holds:
@@
[…]
@@ 1.3. The partial Cartan matrix
[…]
@@ 1.5. Homogeneous Carter diagrams and the transition matrix
[…]
@@ 1.6. Transitions
[…]
@@
[…]
@@ 1.7. Admissible diagrams and Carter diagrams
According to (1.4(a)(a)), there exists a certain set (1.12) of linearly independent roots.Thanks to (1.4(b)(b)), there exists a partition $S=S_{\alpha}\coprod S_{\beta}$ which is said to be the bicolored partition. The admissible diagram is bicolored, i.e., the set of nodes can be partitioned into two disjoint subsets $S_{\alpha}=\{\alpha_{1},\dots,\alpha_{k}\}$ and $S_{\beta}=\{\beta_{1},\dots,\beta_{h}\}$, where roots of $S_{\alpha}$ (resp. $S_{\beta}$) are mutually orthogonal.
@@ 2.1. When a vector is a linkage root?
where $\langle\cdot,\cdot\rangle$ is the usual dot product. Since $B_{\Gamma}$ is positive definite, the eigenvalues of $B_{\Gamma}$ are positive. Therefore, the matrix $B^{-1}_{\Gamma}$ is also positive definite. We call the quadratic form $\mathscr{B}^{\vee}_{\Gamma}$ corresponding to the matrix $B^{-1}_{\Gamma}$ the inverse quadratic form. The form $\mathscr{B}^{\vee}_{\Gamma}$ is positive definite.
@@ 2.2. Criterion that a vector is a linkage root
$^{t}{M}\cdot B_{\Gamma^{\prime}}\cdot M=B_{\Gamma}.\\$
@@ 2.3. Full linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$
[…]
@@
$D_{4}$,  $D_{4}(a_{1})$ 33 $3\times 8=24$ - $24$
$D_{5}$,  $D_{5}(a_{1})$ 33 $10$ $2\times 16=32$ $42$
$D_{6}$, $D_{6}(a_{1})$, $D_{6}(a_{2})$ 33 $12$ $2\times 32=64$ $76$
$D_{7}$, $D_{7}(a_{1})$, $D_{7}(a_{2})$ 33 $14$ $2\times 64=128$ $142$
$D_{l}$,  $D_{l}(a_{k})$, $l>7$ 11 $2l$ - $2l$
@@ 3.1.1. Action of $B^{-1}_{\Gamma}$ on the space $L$
Let $S=\{\tau_{1},\dots,\tau_{l}\}$ be a $\Gamma$-set, $L$ - the space spanned by roots of $S$, $\gamma$ be a root in $L$:
@@ 3.1.2. The projection of the linkage root
Let $L^{\perp}$ be the orthogonal complement of $L$ to $L(\gamma)$ in the sense of the symmetric bilinear form $(\cdot,\cdot)$ associated with the root system $\varPhi$:
@@
(ii) The vector $\mu$ is, up to sign, a fixed vector for any root $\L\theta\in L(\gamma)\backslash$ L.
(iii) The value $\mathscr{B}_{\Gamma}(\theta_{L})$ is constant for any root $\L\theta\in L(\gamma)\backslash$ L.
(iv) The vector $\theta=\theta_{L}+\mu$ is a root if and only if the conjugate vector $\overline{\theta}=\theta_{L}-\mu$ is a root. In addition, $\L\theta\in L(\gamma)\backslash$ L if and only if $\L\overline{\theta}\in L(\overline{\gamma})\backslash$ L.
@@
(vi) If $\delta$ is a root in $\LL(\gamma)\backslash L$ such that $\delta^{\nabla}=\gamma^{\nabla}$, then $\delta=\gamma$ or $\delta=\overline{\gamma}$.
(vii) For any root $\L\theta\in L(\gamma)\backslash$ L, we have
@@
and, $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})$ is a constant for all roots $\L\theta\in L(\gamma)\backslash$ L.
@@
$\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})=\langle B_{\Gamma}^{-1}\theta^{\nabla},\theta^{\nabla}\rangle=\langle{\theta_{L}},B_{\Gamma}\theta_{L}\rangle=\mathscr{B}_{\Gamma}(\theta_{L}).\\$
@@ 3.3. Quadratic forms for homogeneous Carter diagrams
$^{t}{M}\cdot B_{\Gamma^{\prime}}\cdot M=B_{\Gamma},\\$ (3.8)
@@ 3.4. The chain of homogeneous pairs
$\{\{\Gamma_{0},\Gamma_{1}\},\{\Gamma_{1},\Gamma_{2}\},\dots,\{\Gamma_{k-1},\Gamma_{k}\},\text{ where }\Gamma_{0}=\Gamma^{\prime}\text{ and }\Gamma_{k}=\Gamma.\\$ (3.9)
@@
[…]
@@ 3.5. Spectrum of the Dynkin and Carter diagram
Proof. Note that for any transition matrix $M$ and any similarity matrix $L_{\tau}$ the determinant is equal to 1-1, see (1.10). Then, by (3.8) and (3.10), we have
@@
[…]
Therefore, $0\leq\rho\leq 4$. By Proposition 3.1(ii) $B_{\Gamma}$ is positive definite, then $\rho>0$. For the finite Weyl groups, $1$ is not an eigenvalue of the Coxeter transformation, then $\rho\neq 4$. Thus, $0<\rho<4$. ∎
@@ 3.6.2. Action of the Weyl group $W_{S}^{\vee}$
[…]
@@
Proof. Let $\{\tau_{k},\tau_{i}\}$ be a solid edge. If $u^{\nabla}_{\tau_{k}}=u^{\nabla}_{\tau_{i}}=1$ (resp. $u^{\nabla}_{\tau_{k}}=u^{\nabla}_{\tau_{i}}=-1$) then roots $\{-u,\tau_{k},\tau_{i}\}$ (resp. $\{u,\tau_{k},\tau_{i}\}$) constitute the root system corresponding to the extended Dynkin diagram $\widetilde{A}_{2}$, which is impossible, see [St17, Lemma A.1]. For remaining pairs $\{u^{\nabla}_{\tau_{k}},u^{\nabla}_{\tau_{i}}\}$, we have $1\leq u^{\nabla}_{\tau_{k}}+u^{\nabla}_{\tau_{i}}\leq 1$. Now, let $\{\tau_{k},\tau_{i}\}$ be a dotted edge. If $u^{\nabla}_{\tau_{k}}=1$ and $u^{\nabla}_{\tau_{i}}=-1$ (resp. $u^{\nabla}_{\tau_{k}}=-1$ and $u^{\nabla}_{\tau_{i}}=1$) then roots $\{u,-\tau_{k},\tau_{i}\}$ (resp. $\{u,\tau_{k},-\tau_{i}\}$) constitute the root system $\widetilde{A}_{2}$, which is impossible. For remaining pairs $\{u^{\nabla}_{\tau_{k}},u^{\nabla}_{\tau_{i}}\}$, we have $1\leq u^{\nabla}_{\tau_{k}}-u^{\nabla}_{\tau_{i}}\leq 1$. ∎
@@
[…]
@@
[…]
@@ 4.1. The vertex extension
[…]
@@
[…]
@@ 4.2. Partial linkage systems
[…]
@@
[…]
$\gamma_{1}=\mu_{1}+\sum\limits_{j=1}^{l}t_{j}\tau^{1}_{j},\\$ (4.3)
[…]
$\gamma_{2}=\mu_{2}+\sum\limits_{j=1}^{l}t_{j}\tau^{2}_{j},\\$ (4.4)
@@
Since inner products $(\tau^{1}_{j},\tau^{1}_{k})$ and $(\tau^{1}_{j},\tau^{1}_{k})$ correspond to the same edge of $\Gamma^{\prime}$ (or both are equal to 0) we get $\gamma^{\nabla}_{1}=\gamma^{\nabla}_{2}$. Thus, we get